# Demand, and consumer surplus in the payday-loan market: Evidence from British Columbia

### v1

@@ Abstract
This study examines how interest rate caps affect the demand for payday loans, using aggregate data from British Columbia (2012–2019) during which the province’s maximum fee was reduced from \$23 to \$17 and then to \$15 per \$100 borrowed. Estimating a linear demand function via OLS, we find that lowering interest rate caps significantly increases loan demand. We estimate that the \$8 decrease, from \$23 to \$15 per \$100, raised annual consumer surplus by roughly \$28.6 million (2012 CAD). A further reduction to \$14, starting in January 2025, would add another \$3.9 million per year. These results suggest that stricter interest rate caps can yield substantial consumer welfare gains.
@@
Keywords: payday loans; interest rate cap; consumer surplus.
@@ 1 Introduction
Payday loans are typically extended to borrowers with limited liquidity. The high cost of this credit can lead to repeat borrowing and persistent debt cycles (Stegman (2007)). Users of these products in North America are disproportionately young and from marginalized communities; many have low or volatile incomes, precarious employment, and thin or impaired credit histories (Buckland (2018); Apgar and Herbert (2006); Visano (2008)). To reduce the risk of exploitative practices and protect vulnerable consumers, most jurisdictions in the United States and Canada have enacted payday lending regulations. One common measure is a cap on borrowing costs, aimed at curbing the harmful features of payday lending (Kaufman (2013); Dilay and Williams (2018)).
Payday lenders typically price at or near the prevailing cap (Robinson (2018); Fekrazad (2020)). For example, a 15% cap (i.e., a \$15 fee per \$100 borrowed for a two-week loan) translates into an annual percentage rate (APR) of approximately 391%. Where market power exists, price ceilings may enhance efficiency. In addition, interest rate caps can serve as a rudimentary form of social insurance by reallocating surplus from lower-marginal-utility lenders to higher-marginal-utility borrowers (Glaeser and Scheinkman (1998)).
However, direct evidence linking interest rates to the demand for payday loans remains limited. In this study, we exploit policy changes in British Columbia between 2012 and 2019 that reduced the maximum fee per \$100 borrowed from \$23 to \$17, and subsequently to \$15. Using provincial-level aggregate data on borrowing costs and total loan value, we estimate a linear demand curve for payday loans and calculate the associated welfare effects. We estimate that the \$8 reduction from \$23 to \$15 increased consumer surplus by about \$28.6 million per year (2012 CAD). A recent policy effective January 2025 further lowers the cap from \$15 to \$14 per \$100. We estimate that this will add roughly \$3.8 million (2012 CAD) in annual consumer surplus.
@@ 2 Data
Year Borrowing cost Total loan volume Total loan volume
(per CAD 100) (nominal CAD) (2012 CAD)
2012 21.50 318.1 Mil 318.1 Mil
2013 21.70 351.4 Mil 344.4 Mil
2014 21.90 385.3 Mil 370.3 Mil
2015 21.70 340.9 Mil 321.2 Mil
2016 21.70 369.7 Mil 341.5 Mil
2017 19.30 397.3 Mil 359.8 Mil
2018 16.82 416.1 Mil 369.5 Mil
2019 15.30 441.5 Mil 384.4 Mil
2020 14.92 390.7 Mil 333.5 Mil
2021 14.69 271.5 Mil 227.2 Mil
2022 14.85 305.7 Mil 250.8 Mil
2023 14.92 327.2 Mil 263.2 Mil
@@ 3.1 Demand of payday loans
We assume that borrowers face a linear inverse demand curve for payday loans given by:
@@ 3.2 Estimation of demand
Although this yields only eight data points, prior research suggests that even a small sample can produce informative estimates when the underlying relationship is strong and stable (Jenkins and Quintana-Ascencio (2020)). Consistent with theory, the estimated slope coefficient is negative and statistically significant: $\hat{\beta}=-6.61$ million CAD (SE = 2.62 million). This implies that a \$1 increase in the fee per \$100 is associated with about a 6.61 million CAD decrease in annual loan volume. The estimated intercept is $\hat{\alpha}=483.5$ million CAD (SE = 52.7 million). These OLS results are reported in Table 2.
@@ 4 Consumer surplus gains from lower interest rate caps
The CS change from reducing the cap from $p_{1}$ to $p_{2}<p_{1}$ equals the area under demand between $p_{1}$ and $p_{2}$:
$\Delta CS(p_{1}\to p_{2})=\frac{1}{100}\int_{p_{1}}^{p_{2}}\hat{Q}(p)\,dp=\frac{1}{100}\Big[\hat{a}(p_{1}-p_{2})-\frac{\hat{b}}{2}\big(p_{1}^{2}-p_{2}^{2}\big)\Big],$ (2)
@@ 5 Conclusion
Using a small set of annual aggregate data, we estimate a linear demand curve for payday loans and approximate the associated consumer surplus. Despite the coarse granularity of the data—and the necessarily approximate welfare calculations—the results are consistent and policy-relevant. In British Columbia, lowering the cap from \$23 to \$15 per \$100 borrowed coincided with a sizable increase in annual loan volume (from roughly \$331 million to \$384 million, 2012 CAD) and an estimated \$28.6 million gain in consumer surplus per year. A further cap reduction to \$14 per \$100 (implemented nationally in 2025) is projected to add about \$3.9 million annually. These findings suggest that tighter interest-rate ceilings can enhance borrower welfare without suppressing overall loan usage.
The contraction of storefront lenders alongside the rise of online providers points to heterogeneous cost structures: online firms may profitably extend credit at lower fees than brick-and-mortar lenders. Nevertheless, policymakers should account for supply-side adaptation when considering further cap reductions. Much more can be learned from future research using microdata to examine heterogeneity across borrower segments, default risk, and lender margins, and to assess substitution toward other forms of credit. Overall, our evidence supports carefully calibrated caps as a viable consumer-protection tool in high-cost credit markets—provided regulators continue to monitor market composition and credit access as the industry evolves.

### v3

# Demand and consumer surplus in the payday-loan market: Evidence from British Columbia

@@ Abstract
This study examines how interest-rate caps affect the demand for payday loans, using aggregate data from British Columbia from 2012 to 2019, during which the province’s maximum fee was reduced from \$23 to \$17 and then to \$15 per \$100 borrowed. Estimating a linear demand function via OLS, we find that lowering interest-rate caps significantly increases loan demand. We estimate that the \$8 decrease from \$23 to \$15 raised annual consumer surplus by roughly \$28.6 million (2012 CAD). A further reduction to \$14, starting in January 2025, is projected to add an additional \$3.9 million per year. These results suggest that stricter interest-rate caps within a certain range can yield substantial consumer welfare gains, with no evidence of restricted access to payday loans due to lender exit from the market.
@@
Keywords: Consumer finance; Demand; Payday loans; Interest rate cap; Consumer surplus.
@@ 1 Introduction
Payday loans are typically extended to borrowers with limited liquidity. The high cost of this credit can lead to repeat borrowing and persistent debt cycles (stegman2007payday). Users of these products in North America are disproportionately young and from marginalized communities; often women with have low or volatile incomes, precarious employment, and thin or impaired credit histories (buckland2018socio; apgar2006subprime; visano2008different; chen2022national; spotton2018mainstream). To reduce the risk of exploitative practices and protect vulnerable consumers, most jurisdictions in the United States and Canada have enacted payday lending regulations. One common measure is a cap on borrowing costs, aimed at curbing the harmful features of payday lending (kaufman2013payday; dilay2018payday).
Payday lenders typically price at or near the prevailing cap (avery2011payday; robinson2018business). For example, a 15% cap (i.e., a \$15 fee per \$100 borrowed for a two-week loan) translates into an annual percentage rate (APR) of approximately 391%. Where market power exists, price ceilings may enhance efficiency. In addition, interest rate caps can serve as a rudimentary form of social insurance by reallocating surplus from lower-marginal-utility lenders to higher-marginal-utility borrowers (glaeser1998neither).
Studies on interest-rate caps for mainstream consumer loans, which typically have APRs between 6% and 50%, show that increasing interest-rate caps raises the likelihood that credit-constrained consumers can obtain a loan, whereas lowering the cap may restrict access and potentially harm these borrowers (rigbi2013effects; madeira2019impact). However, the population targeted by payday loans is very different from that of mainstream consumer loans, and there is little existing evidence examining how interest rates affect the demand for payday loans and the resulting consumer surplus. Only one earlier study exploited transaction-level data to examine the reduction of the maximum fee per \$100 borrowed from \$15 to \$10 in Rhode Island and estimated an annual consumer-surplus gain of approximately \$3.4 million (USD) (fekrazad2020impacts). In this study, we exploit multiple policy changes in British Columbia—which has a payday loan market roughly five times larger than Rhode Island’s—to analyze cap reductions between 2012 and 2019 that lowered the maximum fee per \$100 borrowed from \$23 to \$17, and subsequently to \$15. Using provincial-level aggregate data on borrowing costs and total loan volume, we estimate a linear demand curve for payday loans and calculate the associated welfare effects. We estimate that the \$8 reduction from \$23 to \$15 increased consumer surplus by approximately \$28.6 million per year (2012 CAD). A more recent policy, effective January 2025, further reduces the cap from \$15 to \$14 per \$100, and we estimate that this change will generate an additional \$3.9 million (2012 CAD) in annual consumer surplus.
@@ 2 Data
It is also worth noting that the number of physical storefront lenders began declining in 2015, rebounded slightly in 2018 and 2019, and has continued to decline since then, while online payday lending in British Columbia began expanding in 2020.
@@
Year Borrowing cost Annual loan volume Total loan volume # of locations # of locations
(per CAD 100) (nominal CAD) (2012 CAD) (H.O. +Branches) (Online only)
2012 21.50 318.1 Mil 318.1 Mil 274 -
2013 21.70 351.4 Mil 344.4 Mil 275 -
2014 21.90 385.3 Mil 370.3 Mil 274 -
2015 21.70 340.9 Mil 321.2 Mil 226 -
2016 21.70 369.7 Mil 341.5 Mil 209 -
2017 19.30 397.3 Mil 359.8 Mil 192 -
2018 16.82 416.1 Mil 369.5 Mil 202 -
2019 15.30 441.5 Mil 384.4 Mil 204 -
2020 14.92 390.7 Mil 333.5 Mil 194 14
2021 14.69 271.5 Mil 227.2 Mil 187 15
2022 14.85 305.7 Mil 250.8 Mil 173 21
2023 14.92 327.2 Mil 263.2 Mil 170 21
@@ 3.1 Demand of payday loans
As in fekrazad2020impacts, we assume that borrowers face a linear inverse demand curve for payday loans, given by:
@@ 3.2 Estimation of demand
Although this yields only eight data points, prior research suggests that even a small sample can produce informative estimates when the underlying relationship is strong and stable (jenkins2020solution). Consistent with theory, the estimated slope coefficient is negative and statistically significant: $\hat{\beta}=-6.61$ million CAD (SE = 2.62 million). This implies that a \$1 increase in the fee per \$100 is associated with about a 6.61 million CAD decrease in annual loan volume. The estimated intercept is $\hat{\alpha}=483.5$ million CAD (SE = 52.7 million). These OLS results are reported in Table 2.
@@ 4 Consumer surplus gains from lower interest rate caps
The CS change from reducing the cap from $p_{1}$ to $p_{2}$ ($p_{2}<p_{1}$) equals the area under demand between $p_{1}$ and $p_{2}$:
$\Delta CS(p_{1}\to p_{2})=\frac{1}{100}\int_{p_{2}}^{p_{1}}\hat{Q}(p)\,dp=\frac{1}{100}\Big[\hat{a}(p_{1}-p_{2})-\frac{\hat{b}}{2}\big(p_{1}^{2}-p_{2}^{2}\big)\Big],$ (2)
@@
5 Robustness Check
To further ensure the reliability of our primary findings, we estimate a quadratic demand curve using 2012–2019 data, modeling real annual payday loan volume $Q(p)$ (in 2012 dollars) as a function of the cost of borrowing $p$ (in dollars per \$100):
$Q(p)=a-b\,p+c\,p^{2},b>0,$
where the estimated coefficients from OLS are:
$\hat{a}=671{,}000{,}000,\quad\hat{b}=26{,}900{,}000,\quad\hat{c}=539{,}434.$
This functional form allows for curvature in borrower response to price, with a slight flattening at higher prices.
Evaluating $\hat{Q}(p)$ at the relevant cap levels yields: $\hat{Q}(23)=337.936$ million, $\hat{Q}(15)=388.872$ million, $\hat{Q}(14)=400.130$ million.
The change in consumer surplus (CS) when price falls from $p_{1}$ to $p_{2}$ is:
$\Delta CS(p_{1}\to p_{2})=\frac{1}{100}\int_{p_{2}}^{p_{1}}\hat{Q}(p)\,dp=\frac{1}{100}\Big[\hat{a}(p_{1}-p_{2})-\frac{\hat{b}}{2}\big(p_{1}^{2}-p_{2}^{2}\big)+\frac{\hat{c}}{3}\big(p_{1}^{3}-p_{2}^{3}\big)\Big].$ (3)
Plugging in $\hat{a}$ and $\hat{b}$, we obtain $\Delta CS_{23\to 15}\approx 28.59$ million (2012 CAD) and $\Delta CS_{15\to 14}\approx 3.94$ million (2012 CAD) per year. These figures are very close to the estimates obtained from the linear demand specification.
@@ 6 Conclusion
Using a small set of annual aggregate data, we estimate a linear demand curve for payday loans and approximate the associated consumer surplus. Despite the coarse granularity of the data—and the necessarily approximate welfare calculations—the results are consistent and policy-relevant. In British Columbia, lowering the cap from \$23 to \$15 per \$100 borrowed coincided with a sizable increase in annual loan volume (from roughly \$331 million to \$384 million, 2012 CAD) and an estimated \$28.6 million gain in consumer surplus per year. A further cap reduction to \$14 per \$100 (implemented nationally in 2025) is projected to add about \$3.9 million annually. These findings are consistent with an earlier study that investigated a similar policy change in the United States, which suggests that tighter interest-rate ceilings can enhance borrower welfare without suppressing overall loan usage (fekrazad2020impacts).
The contraction of storefront lenders alongside the rise of online providers points to heterogeneous cost structures: online firms may profitably extend credit at lower fees than brick-and-mortar lenders. Nevertheless, policymakers should consider potential supply-side responses when evaluating further cap reductions. Prior research suggests that lowering the interest-rate cap from \$15 to \$10 per \$100 borrowed may have contributed to lender exit and reduced consumer access in Oregon (zinman2010restricting). However, these findings should be interpreted with caution, as the average payday-loan default rate in Oregon—approximately 10 to 12 percent—is substantially higher than the 2 to 5 percent range observed in British Columbia. In contrast, a more recent study examining the same policy change in Rhode Island, where default rates are similar to those in British Columbia, found no evidence of reduced access resulting from the cap reduction (fekrazad2020impacts).
Our main findings are based on only a few aggregate data points from a single province in Canada, and much more can be learned from future research using aggregate data from multiple jurisdictions and micro-level data to examine heterogeneity across borrower segments, default risk, and lender margins, as well as to assess substitution toward other forms of credit. Overall, our evidence supports carefully calibrated caps within a certain range as a viable consumer-protection tool in high-cost credit markets—provided that regulators continue to monitor market composition and credit access as the industry evolves.
Funding
No funding was received for this work.
Disclosure statement
No potential conflict of interest was reported by the authors.